OpenStudy (dunk789):

a young lady needs help what x+3=43

OpenStudy (welshfella):

You need to solve for x. To isolate x subtract 3 from each side of the equataion

OpenStudy (dunk789):

so 3-x=0?

OpenStudy (welshfella):

x + 3 = 43 x + 3 - 3 = 43 - 3 now simplify the left side (of the =) and the right side

OpenStudy (dunk789):

what is simplify?

OpenStudy (phi):

most people know how to do 43-3 and make that a number